OpenStudy (anonymous):

Solve by elimination or elimination with multiplication. x - y = 3 and x + y = 7

OpenStudy (anonymous):

1 -1 3 1 1 7 1 -1 3 0 2 4 1 -1 3 0 1 2 1 0 5 0 1 2 x = 5, y = 2

OpenStudy (anonymous):

that makes no sense whatsoever....

OpenStudy (anonymous):

Ok that process seems unclear to me also, but the goal here is to find an x and y for which both equations evaluate. So read this question as "For what value of x and what value of y, gives 3 when you substract them (x-y = 3), and gives 7 when you add them (x+y = 7)? Does that make sense?

OpenStudy (anonymous):

This is a classic 'substitution equation. x - y = 3 so x = y + 3 so substitute x x + y = 7 so y + 3 + y = 7 so 2y = 7 -3 2y =4 y = 4/2 y = 2 x -y = 3 so x - 2 = 3 x = 2 + 3 x =5

OpenStudy (anonymous):

make sense?

OpenStudy (anonymous):

Nikkie that is not right, check again. You're first saying x = 7 and y = 4, which would mean 7-4 = 3 ánd 7 + 4 = 7...

OpenStudy (anonymous):

The x and y values are the same in both formulas. You use different ones.